40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 122 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  10733   Mon Nov 24 20:24:29 2014 diegoUpdateSUSAnti-Jitter Telescope for OpLevs

I stared a bit longer at the plots and thanks to Eric's feedback I noticed I payed too much attention to the comparison between Beta and Gamma and not enough attention to the fact that Beta has some zero-crossings...

I made new plots, focusing on this fact and using some real values for the focal lengths; some of them are still a bit extreme, but I wanted to plot also the zero-crossings for high values of x, to see if they make sense.

 

Plot of Beta and Gamma

 20141124_Plot_Real_BetaGamma_f.pdf

 

 

Plot of Beta and Gamma (zoom)

 

 20141124_Plot_Real_BetaGamma_f_Zoom.pdf

 

If we are not interested in the sign of our signals/noises (apart from knowing what it is), it is maybe more clear to see regions of interest by plotting Beta and Gamma in absolute value:

 

Plot of Beta and Gamma (Abs)

 20141124_Plot_Real_BetaGamma_Abs_f.pdf

 

 

I don't know if putting the telescope far from the QPD and near the mirror has some disadvantage, but that is the region with the most benefit, according to these plots.

 

The plots shown so far only consider the coefficients of the various terms; this makes sense if we want to exploit the zero-crossing of Beta's coefficient and see how things work, but the real noise and signal values also depend on the Alpha and Theta themselves. Therefore I made another kind of plot, where I put the ratio r'(Alpha)/r'(Theta) and called it Tau. This may be, in a very rough way, an estimate of our "S/N" ratio, as Alpha is the tilt of the mirror and Theta is the laser jitter; in order to plot this quantity, I had to introduce the laser parameters r and Theta (taken from the Edmund Optics 1103P datasheet), and also estimate a mean value for Alpha; I used Alpha = 200 urad. In these plots, the contribute of r'(r) is not considered because it doesn't change adding the telescope, and it is overall small.

In these plots the dashed line is the No Telescope case (as there is no variable quantity), and after the general plot I made two zoomed subplots for positive and negative focal lengths.

 

Plot of Tau (may be an estimate of S/N)

20141124_Plot_Real_Tau_f.pdf

 

 

Plot of Tau (positive f)

20141124_Plot_Real_Tau_f_Pos.pdf

 

Plot of Tau (negative f)

20141124_Plot_Real_Tau_f_Neg.pdf

 

If these plot can be trusted as meaningful, they show that for negative focal lengths our tentative "S/N" ratio is always decreasing which, given the plots shown before, it does little sense: although for these negative f Gamma never crosses zero, Beta surely does, so I would expect one singular value each.

Attachment 2: 20141124_Plot_Real_BetaGamma_f_Zoom.pdf
20141124_Plot_Real_BetaGamma_f_Zoom.pdf
Attachment 3: 20141124_Plot_Real_BetaGamma_Abs_f.pdf
20141124_Plot_Real_BetaGamma_Abs_f.pdf
  10732   Fri Nov 21 18:23:01 2014 diegoUpdateSUSAnti-Jitter Telescope for OpLevs

EDIT: some images look bad on the elog, and the notebook is parsed, which is is bad. Almost everything posted here is in the compressed file attachment.

 

As we've been discussing, we want to reduce the laser's jitter effect on the QPDs of the OpLevs, without losing sensitivity to angular motion of the mirror; the current setup is roughly described in this picture:

1.pdf

 

 The idea is to place an additional lens (or lenses) between the mirror and the QPD, as shown in the proposed setup in this picture:

2.pdf

 

 I did some ray tracing calculations to find out how the system would change with the addition of the lens. The step-by-step calculations are done at the several points shown in the pictures, but here I will just summarize. I chose to put the telescope at a variable relative distance x from the QPD, such that x=0 at the QPD, and x=1 at the mirror.

 

Here are the components that I used in the calculations:

 

Propagator

propagator.png

 

Tilted Mirror

tilted_flat_mirror.png

 

Telescope

telescope.png

 

I used a 3x3 matrix formalism in order to have easier calculations and reduce everything to matrix multiplications; that because the tilted mirror has an annoying addictive term, which I could get rid of:

2x2_3x3.png

 

Therefore, n the results the third line is a dummy line and has no meaning.

 

For the first case (first schematic), we have, for the final r and Theta seen at the QPD:

result_old.png

 

 

In the second case, we have a quite heavy output, which depend also on x and f:

 result_new.png

 

Now, some plots to help understand the situation.

What we want if to reduce the angular effect on the laser displacement, without sacrificing the sensitivity on the mirror signal. I defined two quantities:

beta.png

gamma.png

Beta is the laser jitter we want to reduce, while Gamma is the mirror signal we don't want to lose. I plotted both of them as a function of the position x of the new lens, for a range of focal lengths f. I used d1 = d2 = 2m, which should be a realistic value for the 40m's OpLevs.

 

Plot of Beta

20141121_Plot_Real_Beta_f.pdf

 

Plot of Gamma

20141121_Plot_Real_Gamma_f.pdf

 

Even if it is a bit cluttered, it is useful to see both of the same plot:

 

Plot of Beta & Gamma

20141121_Plot_Real_BetaGamma_f.pdf

 

 

 Apart from any kind of horrific mistakes that I may have done in my calculations, it seems that for converging lenses our signal Gamma is always reduced more than the jitter we want to suppress. For diverging lenses, the opposite happens, but we would have to put the lens very near to the mirror, which is somehow not what I would expect. Negative values of Beta and Gamma should mean that the final values at the QPD level are on the opposite side of the axis/center of symmetry of the QPD with respect to their initial position.

 

I will stare at the plots and calculations a bit more, and try to figure out if I missed something  obvious. The Mathematica notebook is attached.

Attachment 14: 141121_antijitter_telescope.tar.bz2
  10731   Fri Nov 21 13:58:51 2014 ericqUpdateOptical LeversHeNe RIN test

 Steve had me measure the RIN of a JDSU HeNe laser. I used a PDA520, and measured the RIN after the laser had been running for about an hour, which let the laser "settle" (I saw the low frequency RIN fall after this period).

Here's the plot and zipped data.

Steve: brand new laser with JDSU 1201 PS

RIN_P893519.pdf

Attachment 1: 2014-11-21_HeNeRIN.zip
  10730   Fri Nov 21 11:41:24 2014 manasaUpdateGeneralX end fiber insulated and on cable tray

Quote:

Quote:

[Steve, Diego, Manasa]

Since the beatnotes have disappeared, I am taking this as a chance to put the FOL setup together hoping it might help us find them.

Two 70m long fibers now run along the length of the Y arm and reach the PSL table.

The fibers are running through armaflex insulating tubes on the cable racks. The excess length ~6m sits in its spool on the top of the PSL table enclosure.

Both the fibers were tested OK using the fiber fault locator. We had to remove the coupled end of the fiber from the mount and put it back in the process. So there is only 8mW of end laser power at the PSL table after this activity as opposed to ~13mW.  This will be recovered with some alignment tweaking.

After the activity I found that the ETMY wouldn't damp. I traced the problem to the ETMY SUS model not running in c1iscey. Restarting the models in c1iscey solved the problem.

 

 AP Armaflex  tube 7/8" ID X 1" wall insulation for the long fiber in wall mounted cable trays installed yesterday.

The 6 ft long sections are not glued. Cable tied into the tray pressed against one an other, so they are air tight. This will allow us adding more fibers later.

 Atm2: Fiber PSL ends  protection added on Friday.

 

[Steve, Manasa] 

Two 70m long fibers are now running through armaflex insulating tubes along the X arm on the cable racks. The excess length of the fiber sits in its spool on top of the PSL enclosure. 

Fibers were checked after this with the fiber fault locator (red laser) and found OK.

  10729   Fri Nov 21 02:22:25 2014 ericqUpdateLSCMore ALS PRFPMI exploration

 Similar to what Jenne did the other night, I kept the PRFPMI arm DoFs locked on ALS, in hopes to check out the RF error signals. 

I was able to stably sit at nominally zero offset in both CARM and DARM, tens of minutes at a time, and the PRMI could reacquire without a fuss. Arm powers would rest between 10 and 20, intermittently exhibiting the "buzzing" behavior that Jenne mentioned when passing through resonance. 100pm CARM offset means arm powers of around 10, so since our ALS RMS is on this order, this seems ok. I saw TRX get as high as 212 counts, which is just about the same that I've simulated as the maximum power in our IFO. 

To get this stable, I turned off all boosts on MICH and PRCL except PRCL FM6, and added matrix elements of 0.25 for TRX and TRY in the trigger line for the PRMI DoFs. The logic for this is that if the arm powers are higher than 1, power recycling is happening, so we want to keep things above the trigger down value of 0.5, even if POP22 momentarily drops. 

I also played around a bit with DARM offsets. We know from experience that the ALS IR resonance finding is not super precise, and thus zero in the CARM FM is not zero CARM offset when on ALS. The same obviously holds for DARM, so I moved the DARM offset around, and could see the relative strengths of flashes change between the arms as expected.

I've written down some GPS times that I'm going to go back and look at, to try to back out some information about our error signals. 

Lastly, there may be something undesirable happening with the TRX QPD; during some buzzing, the signal would fluctuate into negative values and did not resemble the TRY signal as it nominally would. Perhaps the whitening filters are acting up...

  10728   Thu Nov 20 22:43:15 2014 KojiUpdateIOOIMC WFS damping gain adjustment

From the measured OLTF, the dynamics of the damped suspension was inferred by calculating H_damped = H_pend / (1+OLTF).
Here H_pend is a pendulum transfer function. For simplicity, the DC gain of the unity is used. The resonant frequency of the mode
is estimated from the OLTF measurement. Because of inprecise resonant frequency for each mode, calculated damped pendulum
has glitches at the resonant frequency. In fact measurement of the OLTF at the resonant freq was not precise (of course). We can
just ignore this glitchiness (numerically I don't know how to do it particularly when the residual Q is high).

Here is my recommended values to have the residual Q of 3~5 for each mode.

MC1 SUS POS current  75   -> x3   = 225
MC1 SUS PIT current   7.5 -> x2   =  22.5
MC1 SUS YAW current  11   -> x2   =  22
MC1 SUS SD  current 300   -> x2   = 600

MC2 SUS POS current  75   -> x3   = 225
MC2 SUS PIT current  20   -> x0.5 =  10
MC2 SUS YAW current   8   -> x1.5 =  12
MC2 SUS SD  current 300   -> x2   = 600

MC3 SUS POS current  95   -> x3   = 300
MC3 SUS PIT current   9   -> x1.5 =  13.5
MC3 SUS YAW current   6   -> x1.5 =   9
MC3 SUS SD  current 250   -> x3   = 750


This is the current setting in the end.

MC1 SUS POS 150
MC1 SUS PIT  15
MC1 SUS YAW  15
MC1 SUS SD  450

MC2 SUS POS 150
MC2 SUS PIT  10
MC2 SUS YAW  10
MC2 SUS SD  450

MC3 SUS POS 200
MC3 SUS PIT  12
MC3 SUS YAW   8
MC3 SUS SD  500

Attachment 1: MC_OLTF_CLTF.pdf
MC_OLTF_CLTF.pdf MC_OLTF_CLTF.pdf MC_OLTF_CLTF.pdf MC_OLTF_CLTF.pdf MC_OLTF_CLTF.pdf MC_OLTF_CLTF.pdf MC_OLTF_CLTF.pdf MC_OLTF_CLTF.pdf
  10727   Tue Nov 18 22:34:28 2014 JenneUpdateLSCSome other plots from PRFPMI

Quote:

I was able to hold the PRMI on REFL33I&Q, and have ALS CARM and DARM at zero CARM offset.  The arm would "buzz" through the resonance regularly.  I use the word buzz because that's kind of what it sounded like.  This is the noise of the ALS system.

Here is a plot of when the arm powers went pretty high from last night.  CARM and DARM were on ALS comm and diff, PRMI was on REFL33 I&Q.  I set the CARM offset so that I was getting some full arm resonances, and it goes back and forth over the resonance.

The Y axes aren't perfect when I zoom, but the maximum TRX value was 98 in this plot, while the max TRY value was 107. 

MICH_OUT was hitting its digital rails sometimes, and also it looks like PRCL and MICH occasionally lost lock for very brief periods of time. 

Glitch-like events in PRCL_OUT are at the edges of these mini-locklosses. I don't know why POPDC has glitch-y things, but we should see if that's real.

TRXmax98_TRYmax107_CARMdarmOnALS_0carmOffset.png

Okay, I've zoomed in a bit, and have found that, interestingly, I see that both POP22 and POP110 decrease, then increase, then decrease again as we pass through full resonance.  This happens in enough places that I'm pretty sure we're not just going back and forth on one side of the resonance, but that we're actually passing through it.  Q pointed out that maybe our demod phase angle is rotating, so I've made some zoom-in plots to see that that's not a significant effect.  I plot the I and Q phases individually, as well as the total=sqrt(I**2 + Q**2), along with TRY (since the increases and decreases are common to both arms, as seen in the plot above).

For POP 22:

Zoom_with_POP22.png

For POP 110:

Zoom_with_POP110.png

I also plot the MICH and PRCL error signals along with TRY and POP22 total.  You can see that both MICH and PRCL were triggered off about 0.1msec after POP22 total this it's first super low point.  Then, as soon as POP22 becomes large enough, they're triggered back on, which happens about 1.5msec later. (The triggering was actually on POP22I, not POP22tot, but the shapes are the same, and I didn't want to overcrowd my plots).

Zoom_with_ErrSigs.png

I am not sure if we consistently lose sideband signal in the PRC more on one side of the CARM resonance than the other, but at least POP22 and POP110 are both lower on the farther side of this particular peak.  I want to think about this more in relation to the simulations that we've done.  One of the more recent things that I see from Q is from September:  elog 10502, although this is looking specifically at the REFL signals at 3f, not the 2f circulating PRCL power as a function of CARM offset.

  10726   Tue Nov 18 17:11:30 2014 JenneUpdateLSCSome lockloss plots from PRFPMI

I am still staring at / trying to figure out the latter 4 locklosses posted earlier.  But, I have just included the transmission QPD angular output signals to the frames, so we should be able to look at that with locklosses tonight. 

To get the lockloss plots:  in ..../scripts/LSC/LocklossData/ , first run ./FindLockloss.sh <gps time> .  This just pulls the TRX and TRY data, and doesn't save it, so it is pretty quick.  Adjust the gps time until you capture the lockloss in your plot window.  Then run ./LockLossAutoPlot.sh <gps time> to download and save the data.  Since it has become so many channels, it first makes a plot with all of the error and control signals, and then it makes a plot with the power levels and angular signals.  The data folder is just called <gps time>.  I have started also including a text file called notes inside of the folder, with things that I notice in the moment, when I lose lock.  Don't use .txt for the suffix of the notes file, since the ./PlotLockloss.py <folder name> script that will plot data after the fact tries to plot all .txt files.  I have also been appending the folder name with keywords, particularly _notInteresting or _unknown for either obvious lockloss causes or mysterious lockloss cases.

 

  10725   Tue Nov 18 15:20:58 2014 JenneUpdateLSCSome lockloss plots from PRFPMI

Elog from ~5am last night:

Tonight was just several trials of PRFPMI locking, while trying to pay more attention to the lockloss plots each time.

General notes: 

I tried once to acquire DRMI on 1f while the arms were held off resonance.  I wasn't catching lock, so I went back to PRMI+arms.  I aligned the PMC, which I noted in a separate elog.  

I was able to hold the PRMI on REFL33I&Q, and have ALS CARM and DARM at zero CARM offset.  The arm would "buzz" through the resonance regularly.  I use the word buzz because that's kind of what it sounded like.  This is the noise of the ALS system.

I think we want to add the transmission QPD angular signals to the frames.  Right now, we just keep the sums.  It would have been handy to glance at them, and see if they were wiggling in the same way that some other signal was waggling. 

All the data files are in /opt/rtcds/caltech/c1/scripts/LSC/LocklossData.  Each folder is one lockloss.  It includes text files for each trace, as well as any plots that I've made, and any notes taken.  The text files are several MB each, so I'm not going to bog the elog down with them.  There are a few folders that end in "_notInteresting".  These ones are, as one might guess, not interesting.  2 were MC locklosses (I'm not actuating on MC2, so I declared these independent from my work) and one was when I knew that my ALS was bad - the beatnotes weren't in good places, and so the ALS noise was high.


 Folder:  1100342268_POP22goesLow

Working notes:  Lost lock because POP22 went too low.  PRCL and MICH triggered off.  After this, changed PRCL and MICH "down" thresholds to 0.5, from 10.
 

Plots:

ErrorSignals_NothingObviouslyOscillating.pngErrorSignals_Zoom_MICHprclWeird.pngPowers_POP22goesLow.png

Conclusion:  Easy fix.  Changed the down thresholds for MICH and PRCL to be lower, although still low enough that they will trigger off for a true lockloss.  Why though do we lose so much sideband power when the arm transmission goes high?  POP22 dipped below 10 when TRX went above 29.  Does this happen on both sides of the CARM offset?  Quick simulation needed.

------------------------------------------------------------------------

Folder:  1100330534_maybePRCLangular

Working notes:  PRFPMI, reducing CARM offset to arm powers of 7.  CARM on sqrtInv, DARM on DCtrans. PRMI on REFL33 I&Q. Don't know why I lost lock.  Maybe angular stuff in PRC? I think POP spot was moving in yaw as it started to go bad.
Note, later:  regathered data to also get POP angular stuff. Don't think it's POP angular.  Not sure what it is.

Plots:

ErrSigs_NothingJumpingAtMe.pngNotPOPangular.png

Conclusion:  I'm not sure what this lockloss was caused by, although it is not something that I can see in the POP QPD (which was my initial suspicion).  It is, like many of the rest of the cases, one where I see significant bounce and roll mode oscillations (error and control signals oscillating at 16 and 24 Hz).  I don't think those are causing the locklosses though.

--------------------------------------------------------------------

Folder:  1100334680_unknown_highArmPowers

Working notes:  PRFPMI, carm_up script finished, sitting at arm powers of 8.  CARM, DARM on DC trans.  PRMI on REFL33.   Don't know why lost lock.

Plots:

[Don't have any? - I'll make some]

Conclusion:  Again, I see 16 and 24 Hz oscillations, but I don't think those are causing the lockloss.

---------------------------------------------------------------------

Folder: 1100331950_unknown

Working notes: PRFPMI, arms about 8.  CARM, DARM on DC trans.  PRMI on REFL33.  Don't know why I lost lock.

Plots:

More16and24Hz.png

Conclusion: Don't have an answer.

--------------------------------------------------------

Folder: 1100345981_unknown

Working notes: Lockloss while going to arm powers of 7ish from 6ish.  Not POP angular, POP22 didn't go low.

Plots:

ErrSigs_16and24Hz.pngNotPOPsFault.png

Conclusions:  This one wasn't from POP22 going too low, but again, I don't see anything other than 16 and 24Hz stuff.

 

  10724   Mon Nov 17 23:04:51 2014 JenneUpdatePSLAligned PMC

I aligned the beam into the PMC, mostly in yaw.  Don't know why it drifted, but it was annoying me, so I fixed it.

  10723   Mon Nov 17 20:40:29 2014 rana, diegoUpdateIOOInvestigating the IMC WFS situation

We've known for years that the IMC WFS sensing chain is pointlessly bad, but until recently, we haven't thought it was worth it to fix.

There are problems in all parts of the chain:

  1. The WFS Photodetectors oscillate ~200 MHz when turned up to full gain. Diego and I confirmed this today by measuring the RF spectrum of the signals going into the WFS demod boards and seeing the oscillation change (not much) with RF gain. I recommend we switch the heads into the full gain mode (turn all of the attenuators OFF). At the moment we are operating with the 2dB and 8dB attenuators ON.
  2. The demod board has some bad gain allocation and noisy opamps.
  3. The whitening board has too much up/down of gain with noise injection along the way. And the range cannot fill up the ADC.
  10722   Mon Nov 17 20:28:17 2014 ranaSummaryIOOMC servo summing amp

I modified the /cvs/cds/caltech/target/c1psl/psl.db file to adjust the records for the FSS-FAST signal (to make it go yellow / red at the correct voltages). This was needed to match 5V offset which Koji added to the output of the FSS board back in August.

I also manually adjusted the alarm levels with caput so that we don't have to reboot c1psl. Beware of potential tiimebomb / boot issues if I made a typo! psl.db update in the SVN (also, there were ~12 uncomitted changes in that directory....please be responsible and commit ALL changes you make in the scripts directory, even if its just a little thing and you are too cool for SVN)

  10721   Sat Nov 15 21:51:09 2014 ranaUpdatePEMSeismometers set up for huddle test

  In order to do high quality huddle subtraction, we need to align the seismometer axes to high precision. We would need 1000x subtraction to see the instrument noise floor, but are likely to only get 100x. For that we need to align the axes to 0.5 deg (or do a Wiener coordinate transform with the data). To do this, we need to use a high quality bubble level and eventually iterate after trying out.

We should strain relieve the seismometer cables on the slab. It should be a tight clamp so that acoustic vibrations on the cables are terminated at the clamp and don't get to the seismometers. The clamp can be attached to the slab using some strong epoxy.

  10720   Fri Nov 14 20:31:13 2014 ericqUpdateLSCRIN in transmission a problem?

I took a quick look at single arm RIN. Actuating on MC2 vs. the ETM, or using AS55 instead of POY11 made no noticeable difference in the arm cavity RIN.  Not too surprising, but there it is.

  10719   Fri Nov 14 19:28:33 2014 JenneUpdateSUSASS gain reduced for Yarm

[Koji, Jenne]

We noticed last night that the yarm couldn't handle the old nominal gain for the ASS servos.  We were able to run the ASS using about 0.3 in the overall gain.  So, I have reduced the gain in each of the individual servos by a factor of 3, so that the scripts work, and can set the overall gain to 1.

  10718   Fri Nov 14 17:08:17 2014 JenneUpdateLSCRIN vs. Seismic

T-240 has a different convention than we use.  It assumes that North is aligned with the Y-axis.  Since this is the new guy, and we've been using the Guralps with North = X for many years, Diego and I rotated the T-240, and put a label on it that N/S is Y, and E/W is X.  Obviously Vert is still Z.

  10717   Fri Nov 14 15:45:34 2014 JenneUpdateCDSComputers back up after reboot

[Jenne, Q]

Everything seems to be back up and running.

The computers weren't such a big problem (or at least didn't seem to be).  I turned off the watchdogs, and remotely rebooted all of the computers (except for c1lsc, which Manasa already had gotten working).  After this, I also ssh-ed to c1lsc and restarted all of the models, since half of them froze or something while the other computers were being power cycled. 

However, this power cycling somehow completely screwed up the vertex suspensions.  The MC suspensions were fine, and SRM was fine, but the ITMs, BS and PRM were not damping.  To get them to kind of damp rather than ring up, we had to flip the signs on the pos and pit gains.  Also, we were a little suspicious of potential channel-hopping, since touching one optic was occasionally time-coincident with another optic ringing up.  So, no hard evidence on the channel hopping, but suspicions.

Anyhow, at some point I was concerned about the suspension slow computer, since the watchdogs weren't tripping even though the osem sensor rmses were well over the thresholds, so I keyed that crate.  After this, the watchdogs tripped as expected when we enabled damping but the RMS was higher than the threshold.

I eventually remotely rebooted c1sus again.  This totally fixed everything.  We put all of the local damping gains back to the values that we found them (in particular, undoing our sign flips), and everything seems good again.  I don't know what happened, but we're back online now. 

Q notes that the bounce mode for at least ITMX (haven't checked the others) is rung up.  We should check if it is starting to go down in a few hours.

Also, the FSS slow servo was not running, we restarted it on op340m.

  10716   Fri Nov 14 15:26:45 2014 SteveUpdatesafetyDiego gets safety traning

Diego Bersanetti received 40m specific safety training today.

  10715   Fri Nov 14 11:07:59 2014 manasaUpdateCDSNot able to run models on FE machines

Quote:

PRM, SRM and the ENDs are kicking up.  Computers are down.  PMC slider is stuck at low voltage.

Still not able to resolve the issue.

Except for c1lsc, the models are not running on any of the FE machines . I can ssh into all the machines but could not restart the models on the FE using the usual rtcds restart <modelname>

Something happened around 4AM (inferring from the Striptool on the wall) and the models have not been running since then.

 

Attachment 1: CDSissue.png
CDSissue.png
  10714   Fri Nov 14 08:25:29 2014 SteveUpdateSUSmorning issues

PRM, SRM and the ENDs are kicking up.  Computers are down.  PMC slider is stuck at low voltage.

Attachment 1: morning.png
morning.png
  10713   Fri Nov 14 02:43:05 2014 ericqUpdateLSCPRFPMI HOM resonances

I've extended my analysis to the PRFPMI case, with the current working knowledge of radii of curvature and cavity lengths. However, losses were not included.

I do not see any HOM activity within about 20nm of the carrier TM00 resonance. 

Basically, what I did was use the standard formulae for the reflection and transmission coefficients of FB cavities viewed as compound mirrors. However, I modified the normal spatial propagation terms to include the additional Guoy phase accumulated by the HOMs. I created these coefficients for each arm individually, and then used (rX + rY)/2 as a mirror in the PRC, and used that to create the transmission coefficient for the PRFPMI as a whole, as a function of frequency offset from the carrier, spatial mode order and CARM offset. As a check, this produced the correct finesse for the carrier lock to the single arm and PRFPMI. 

Here is a PRFPMI CARM FSR of all of the fields' power transmission coefficients, up to order n+m=5. 

HOMcurves.pdf 

One can observe some split peaks. There are two causes, the biggest effect is the mismatch between ETM radii of curvatures (ETMX:59.48, ETMY:60.26):, followed by asymmetric arm length(X:37.79, Y:37.81). (I judged this by the visual change of the plot when changing different factors). 

In the following plot, I broke down the peaks by mode order:

 HOMpeaks.pdf

Code, plots attached!

 

Attachment 3: prfpmiHOM.zip
  10712   Thu Nov 13 23:42:01 2014 JenneUpdateLSCRIN vs. Seismic

After Kate, Diego and I moved the seismometers to the corner for a huddle test (see elog 10711), I wanted to check the coherence between the seismometers and the arm transmissions.  

First of all, it looks like either the Guralp or the T-240 have their X and Y backwards. Diego is going to check this tomorrow. 

Between 0.9Hz - 3.5Hz, we have pretty strong coherence with the horizontal seismic channels.  Interestingly, between 8-10Hz, the Yarm has pretty strong coherence with the Z-axes of the seismometers (the coherence is only about 0.6, but it's consistent over a 2-ish Hz wide band). 

The MC transmission doesn't have as much coherence with the seismic, which surprised me.  So, we can try some FF to the MC, but we may also have to do some to the arms.

Seismic_TRXTRYandMC_13Nov2014.pdf

  10711   Thu Nov 13 22:52:48 2014 kateUpdatePEMSeismometers set up for huddle test

Jenne, Diego, Kate

We want to conduct a huddle test with the three 40m seismometers (2 Guralps and 1 Trillium), so we began to get that set up in order. All three are currently sitting on the large granite slab approximately halfway down the length of the MC tube. We had to move all three seismometers: the Trillium had been next to the BS and the Guralps at the end stations. All three are balanced and aligned and we have put the foam box over them. 

The Trillium has not yet been used here, so we had to first wire its power supply. We're now providing its readout box with +/- 20 V. Getting that hooked up required powering down several electronics racks, which involved auxiliary prep work like turning off the suspension watchdogs. We also installed 3 new BNC cables to carry the Trillium x,y,z signals from its box to the CDS AA board. We're using the inputs which had previously been used for recording the STS2 signals. 

We could find only one of the two 'short' Guralp cables, so at the moment just one of the two Guralps is powered and connected to CDS. Jenne made (some time ago) new cables so that we could leave the long cables that run from the corner to each end station in place to preserve the nominal setup. 

Attachment/edit by Jenne: Seismic spectra.  Note that the T240 is connected to the channels that are called STS_1.  I compared the Guralp spectra to our seismic_ref, and they match up pretty well.  The new spectra is maybe a factor of 2 or so above the reference, at a few Hz. Anyhow, the Guralp seems fine.  I am sure that somewhere we have a second short (as in, not 50m long) Guralp cable, I just can't remember right now where it might be.  Also, the T-240 has some seriously crazy noise up around 30Hz.  What's up with that??  I want to ask Zach if he saw this when he had the Trillium, or if it is new.

Seismic_13Nov2014.pdf

Attachment 2: seism_closed.jpg
seism_closed.jpg
Attachment 3: seism_open.jpg
seism_open.jpg
  10709   Thu Nov 13 04:28:28 2014 JenneUpdateLSCRIN in transmission a problem?

[Jenne, Rana, Koji]

We did some thinking on what could be causing the excess RIN that we see in the arm transmissions but not in the MC transmission.  Unfortunately, I don't think we have anything conclusive yet. 

We thought about:

  • Test mass oplevs
  • Input tip tilt jitter
  • MC motion

Oplevs

As Rana reported in elog 10708, we tuned the oplev servos for ITMX, ETMX, ITMY and ETMY.  They all now look like the new SRM oplevs that Rana described in elog 10694.  However, when we re-looked at the RIN after the oplev tuning, we did not see a noticeable change.  So, fixing up the oplevs didn't fix up the RIN.

Side notes:

  • Several optics had low gains for suspos, which were increased to give Qs of ~5ish.
  • When we gave ITMX a 500 count step in pitch (the same size used for all other optics in both pit and yaw), it didn't come back afterward.  This is a little disconcerting.  Rana had to move the alignment slider to get it back so that we had MICH fringing at the AS port again.

Input Tip Tilts

Koji did some work, reported in elog 10706, on how much the jitter of the input pointing tip tilts should affect us.  We don't think that they are moving enough to be the cause of the excess RIN that we see.


Mode Cleaner Motion

We see some coherence between MC2 suspit and TRX/TRY, so we suspect that the MC's motion could be causing problems. 

I looked at the WFS vs. TRX & TRY, and saw significant coherence at the 3 Hz stack resonance.  I think it's clear that the WFS can help suppress this motion more.  The WFS noise level was too bad to see any other coherence at other frequencies. (Side note:  We should consider increasing the analog WFS signal.  As Rana mentioned back in 2008 in elog 454, the signal is super small.  Increasing the analog gain could allow us to suppress motion at more frequencies, although it would be a bit of a pain.)

To try and get some more signal, I routed the IPPOS beam over to the PRM oplev temporarily.  The idea was to be able to look at the IPPOS port, but with a fast channel.  I turned off the BS/PRM HeNe, and removed the last steering mirror before the QPD.  I put in 2 Y1 steering mirrors to get the IPPOS beam across the table and pointing at IPPOS.  I took my measurements 3 times, with different beam sizes on the QPD, to try to image different gouy phases.  I used absorptive ND filters (0.6 + 0.1) to get the light level on the PD such that I had about 10,000 counts per quadrant, where 16,000 counts seemed to be the saturation point. I also reset the dark offsets of the QPD quadrants, although they were so small that I don't think it did much.  I also took out the optical lever calibration from counts to microradians, since that number isn't meaningful for what I was doing.  So, the IPPOS signals (using the PRM oplev channels) are in raw ADC counts.  The first measurement had no lens, and the beam was probably at least half the size of the QPD.  The second measurement had a lens, and the beam on the QPD was about half the original size.  The third measurement had the lens closer to the QPD, so that the beam was about 1mm on the diode.  In none of these cases do I see any significant coherence with the TRX/TRY RIN signals, except at 3 Hz. After my measurements I put the oplev back including all of the digital settings, although for now I just left the IPPOS beam dumped on a razor dump, since it wasn't being used anyway.  I need to realign IPPOS when it's not the middle of the night.

Some thoughts that we have:

  • Maybe it's time to resurrect seismic feedforward on MC length, to suppress some of this 3 Hz motion?
  • Maybe we should be using the MC_L path to offload some of the frequency feedback, so that we're not pushing on MC2 so hard (because if we have bad F2P coupling, this creates beam motion)

I have plots for each of my IPPOS vs. TRX/TRY measurements.  The data is attached.  For each, I looked at the transfer function between IPPOS (using the SUS-PRM_OPLEV channels) and TRX/TRY to get the 'calibration' between input beam motion and transmission RIN.  In all cases, at 3 Hz the coefficient was about 1, so in the power spectra on the right side, there is no calibration applied to the IPPOS signals. 

IPPOS vs. Transmission RIN, no lens, big beam on QPD:

RIN_TRX_TRY_IPPOS_NoLens_12Nov2014.pdf

(Just kidding about the other 2 plots - the elog can't handle it.  They're in the zippyzip file though, and I don't think they look qualitatively different from the no-lens case).

 

Attachment 1: zippyzip.zip
  10708   Thu Nov 13 01:03:28 2014 rana, jenneUpdateSUSOL updates on ITMs and ETMs

 We copied the new SRM filters over onto the OL banks for the ITMs and ETMs. We then adjusted the gain to be 3x lower than the gain at which it has a high frequency oscillation. This is the same recipe used for the SRM OL tuning.

Before this tune up, we also set the damping gains of the 4 arm cavity mirrors to give step response Q's of ~5 for all DOF and ~7-10 for SIDE.

  10707   Thu Nov 13 01:00:37 2014 ranaUpdateLSCRIN in transmission a problem?

 

 I modified the MC2 trans optical setup a little bit: the reflection from the QPD was not dumped and so it was hitting the wall of the black box.

I angled the QPD slightly and moved the dump so that the reflection hit it. The leakage through the 50/50 steering mirror for the QPD was already being dumped and I made sure that that one stayed dumped on its razor dump. After doing this we turned off the WFS and re-aligned the MC2 trans beam onto the QPD to zero the pit/yaw signals. 

  10706   Wed Nov 12 22:22:11 2014 KojiSummaryIOOEstimation of the angular jitter imposed by the TTs

[Koji, Rana, Jenne]

One coil of the TT produce 36nrad/rtHz at DC.

- C1:IOO-TT2_UL_EXC was excited with 5 count_pk at 0.04Hz.

- LSC_TRY exhibited the symmetric reduction of the transmission from 0.95 to 0.90.

1 - (theta/theta0)^2 /2 = 0.90 / 0.95

=> theta / theta0 = 0.32

- 40m beam waist radius is 3.1mm. This means the divergence angle is 1.1e-4 rad.

=> 1.1e-4*0.32 = 3.6e-5 rad

=> 3.6e-5/5 = 7.2 urad/count (per coil)

- DAC noise 1/sqrt(12 fs), where fs is the sampling rate (fs = 16384)

=> 0.002 cnt/rtHz

- One coil causes 7.2u*0.002 = 14 nrad/rtHz (at DC)

- One suspension cause 29 nrad/rtHz (at DC)

Attachment 1: 03.png
03.png
  10705   Wed Nov 12 21:18:32 2014 ericqUpdateLSCDRFPMI, PRFPMI HOM resonances

So, with my last entry, I was guilty of just throwing stuff into the simulation and not thinking about physics... so I retreated to Siegman for some algebraic calculations of the additional Guoy phase accumulated by the HOMs in the arms -> their resonant frequencies -> the arm length offset where they should resonate. Really, this isn't completely precise, as I treated the arms independently, with slightly differing ETM radii of curvature, but I would expect the "CARM Arm" to behave as a sort of average of the two arm cavities in this regard. (EDIT: Also, I didn't really consider the effect of the coupled vertex cavities... so there's more to be done)

The basic idea I used was:

  • Assume ITMs are effectively flat, infinite Rc
  • Use 40mwiki values for ETM curvatures
  • Each additional HG order adds arccos(sqrt(1 - Larm/Rc)) of Guoy phase for a one way trip down the cavity (Eqn 19.19 in Sigman)
  • For each HOM order up to 5 of the carrier and first order sidebands, add the appropriate phase shift 
  • fold it onto +-FSR/2 of the carrier 00 resonance, convert to m

In practice, I threw together a python script to do this all and print out a table. I've highlighted the values within 10nm, but the closet one is 3.8nm

Results:

########## X Arm HOM Resonance Locations in nm ##########
Mode Order:      0     ,      1     ,      2     ,      3     ,      4     ,      5     

Carrier   :          +0,     +156.21,     -219.58,     -63.376,     +92.832,     +249.04
LSB 11    :     +59.563,     +215.77,     -160.02,     -3.8126,      +152.4,      -223.4
USB 11    :     -59.563,     +96.645,     +252.85,     -122.94,     +33.269,     +189.48
LSB 55    :     -234.18,     -77.975,     +78.233,     +234.44,     -141.35,     +14.857
USB 55    :     +234.18,     -141.61,       +14.6,     +170.81,     -204.98,     -48.776


########## Y Arm HOM Resonance Locations in nm ##########
Mode Order:      0     ,      1     ,      2     ,      3     ,      4     ,      5     

Carrier   :          +0,     +154.82,     -222.35,     -67.531,     +87.292,     +242.11
LSB 11    :     +59.313,     +214.14,     -163.04,      -8.218,      +146.6,     -230.57
USB 11    :     -59.313,      +95.51,     +250.33,     -126.84,     +27.978,      +182.8
LSB 55    :     -235.43,     -80.611,     +74.212,     +229.04,     -148.14,     +6.6809
USB 55    :     +235.43,     -141.74,      +13.08,      +167.9,     -209.27,     -54.452
 
Code is attached. Hopefully no glaring mistakes!
 
 
Attachment 1: HOMlist.py.zip
  10704   Wed Nov 12 20:11:41 2014 KojiUpdateIOOMC WFS gain reduced again

MC WFS was oscillative at 1Hz. I've reduced the servo gain further (x1, x1, x10, x1, x1, and x10).

The MC mirrors were realigned, and the WFS offsets were reset.

  10703   Wed Nov 12 18:08:35 2014 JenneUpdateLSCRIN in transmission a problem?

In my previous meditations about RIN, particularly elog 10258, I was only thinking about the RIN contribution at the offset that I was currently sitting at.  Also, In elog 10258 I was comparing to the ALS signals and just said that the trans signals are better which is true, although isn't super helpful when thinking of reduced CARM offsets. 

My summary today is that I think we want to reduce the RIN in arm transmissions by a factor of 3.


Rather than dig around, I just remeasured the RIN, for both the single arm transmissions and the MC transmission.  (Data attached as .xml file)

The RMS RIN for the Xarm is 1.3e-2.  The RMS RIN for the Yarm is 8.9e-3.  The RMS RIN for MCtrans is 4.0e-3.  For the simulations below, I will use 1e-2 as an average RIN for the arms.

RIN_TRX_TRY_MCtrans_12Nov2014.pdf


As an estimate of the RIN's contribution to cavity fluctuations, I divide the RIN by the slope of the CARM transmission peak.  The slope (from optickle) gives me [ delta-W / delta-m ], and the inverse of that gives me [ delta-m / delta-W ].  I multiply this by RIN, which is [ delta-W / W ] to get [delta-m / W]. 

Then, since I'm using the DC transmission signals as my error signals, I use just TRX (normalized to be 1 for single arm resonance) as my Watts.

So, in total, the traces plotted are { TRX * RIN / slope }. 

The 2 plots are the same data, one with linear-x and the other with log-x.  They both include my estimate of the cavity length fluctuations due to RIN at the arm transmission, as well as an estimate of the cavity length fluctuations if the arm RIN was as good as the MC RIN.  I also show the DRFPMI CARM linewidth (23 pm for HWHM), and 1% of that linewidth.  The last trace is 1% of the half-width of the transmission peak, at the current CARM offset.  For example, 1000 pm away from full resonance the half-width is 1000 pm and 1% of that is 10 pm. 

 RINcontribution_12Nov2014_linearXscale.pngRINcontribution_12NOv2014_logXscale.png

What we want to see here is that the solid blue line is below one of the dotted lines.  I think that using the overall linewidth (purple dotted line) isn't really the right thing to look at.  Our goal is to prevent excursions that will get too close to the resonance peak, and cause a lockloss.  A one picometer excursion is a much bigger problem (relatively) below say 100 pm, as opposed to above 100 pm.  So, I think that we should be looking at the half-width of the resonance peak at whatever the current CARM offset is (orange dotted line).  Above 25 pm, the blue line is below the orange line for all offsets plotted.  If we made the arm RIN as good as the MC RIN, that would be true down to 12-ish pm. 

We should be able to safely transition to non-normalized RF signals at 10pm or below.  This implies that (since any RF signals normalized by this RIN-y trans signal will have the RIN), we want to improve the RIN of the transmission PDs by about a factor of 3. (This will lower the blue line such that it crosses the orange dotted line at 10 pm).

 

Attachment 1: RIN_TRX_TRY_MCtrans_12Nov2014_zip.xml.gz
  10702   Wed Nov 12 15:14:09 2014 manasaUpdateGeneralX end AUX laser fiber setup

 I looked at the endtable for possible space to setup optics in order to couple the X end laser into a PM fiber.

Attached is the layout of where the setup will go and what are the existing stuff that will be moved.

ETMXtable.png

  10701   Wed Nov 12 03:22:23 2014 JenneUpdateLSC3f DRMI sensing mat

Koji pointed out something to me that I think he had told me ages ago, and Rana alluded to last night:  Since I'm not tuning my "demod phase" for the sensing matrix lockins, unless I happened to get very lucky, I was throwing away most of the signal.  Lame.  

So, now the magnitude is sqrt(real^2 + imag^2), where real and imag here are the I and Q outputs of the lockin demodulator, after the 0.1Hz lowpass.  (I put in the low pass into all of the Q filter banks).  To keep the signs consistent, I did do a rough tuning of those angles, so that I can use the sign of the real part as the sign of my signal.  When I was PRMI locked, I set the phase for all things acutated by MICH to be 79deg, all things actuated by PRCL to be 20 deg, and when DRMI locked set all things SRCL to be 50deg. 

After doing this, the phases of my sensing matrix output matches Koji's careful analyses.  I don't know where the W/ct numbers I was using came from (I don't think I made them up out of the blue, but I didn't document where they're from, so I need to remeasure them).  Anyhow, for now I have 1's in the calibration screen for the W/ct calibration for all PDs, so my sensing matrices are coming out in cts/m, which is the same unit that Koji's analysis is in. (Plot for comparing to Koji is at end of entry).

While reducing the CARM offset, I left the sensing matrix lines on, and watched how they evolved.  The phases don't seem to change all that much, but the magnitudes start to decrease as I increase the arm power.

For this screenshot, the left plot is the phases of the sensing matrix elements (all the REFL signals, MICH and PRCL), and the right plot is the magnitudes of those same elements.  Also plotted is the TRX power, as a proxy for CARM offset.  The y-scale for the TRX trace is 0-15.  The y-scale for all the phases is -360 to +360.  The y-scale of the magnitude traces are each one decade, on a log scale.

SensMatVsPower_UpToArms10.png

Bonus plot, same situation, but the next lock held for 20 minutes at arm powers of 8.  We don't know why we lost lock (none of the loops were oscillating, that I could see in the lockloss plot).

PRMI_arms8_20minutes.png


Here are some individual sensing matrix plots, for a single lock stretch, at various times.  One thing that you can see in the striptool screenshots that I don't know yet how to deal with for the radar plots is the error bars when the phase flips around by 360 degrees.  Anyhow, the errors in the phases certainly aren't as big as the error boxes make them look.

PRMI just locked, CARM offset about 3nm, CARM and DARM on ALS comm and diff, arm powers below 1:

SensMatMeas_11Nov2014_PRMIarms_ArmPowSmall.png

PRMI still on REFL33 I&Q, CARM and DARM both on DC transmissions, arm powers about 4:

SensMatMeas_11Nov2014_PRMIarms_ArmPow3pt8.png

CARM offset reduced further, arm powers about 6:

SensMatMeas_11Nov2014_PRMIarms_ArmPow6.png

CARM offset reduced even more, arm powers about 7:

SensMatMeas_11Nov2014_PRMIarms_ArmPow7.png


For this plot for comparing with Koji's analysis, I had not yet put 1's in the calibration screen, so this is still in "W"/m, where "W" is meant to indicate that I don't really know the calibration at all.  What is good to see though is that the angles agree very well with Koji's analysis, even though he was analyzing data from yesterday, and this data was taken today.  This sensing matrix is DRMI-only (no arms), 1f locking.

SensMatMeas_11Nov2014_DRMI_fixedMags.png

Bonus plot, PRMI-only sensing matrix, with PRMI held using REFL 33 I&Q:

SensMatMeas_11Nov2014_PRMI_fixedMags.png

 

  10700   Wed Nov 12 01:30:39 2014 ericqUpdateLSCDRFPMI, PRFPMI HOM resonances

I did some simulations to see if we are susceptible to HOM resonances as we reduce the CARM offset. I restricted my search to HG modes of the Carrier+[-55,-11,0,+11,+55]MHz fields with n+m<6, and used all the real physical parameters I could get ahold of. 

In short, as I change the CARM offset, I don't see any stray resonances within 2nm of zero, either in PRFPMI or DRFPMI. 

Now, the mode matching in my simulation is not the real mode matching our real interferometer has. Thus, it can't tell us how much power we may see in a given mode, but it can tell us about our susceptibility to different modes. I.e. if we were to have some power in a certain mode coming out of the IMC, or present in the vertex, we can see what it would do in the arms. 

Since my simulation has some random amounts of power in each HOM coming into the interferometer, I simply swept the CARM offset and looked for peaks in the power of each mode. Many of the fields exhibited gentle slopes over the range, and we know we ok from 3nm->~100pm, so I made the selection rule that a "peak" must be at least 10 times as big as the minimum value over the whole range, in order to see fields that really do have CARM dependence. 

In the following plots, normalized IFO power is plotted and the locations of HOM peaks are indicated with circles; their actual heights are arbitrary, since I don't know our real mode content. However, I'm not really too concerned, since all I see is some -11MHz modes between 2-3nm of full resonance, where we have no problem controlling things... Also, all of the carrier HOMs effectively co-resonate with the 00 mode, which isn't too surprising, and I didn't include these modes in the plots.

 DRFPMI_HOMscan.pngPRFPMI_HOMscan.png

Finally, I visually inspected the traces for all of the modes, and didn't really find anything else peeking out. 

Code, plots attached.  

Attachment 1: HOMs.zip
  10699   Wed Nov 12 00:55:56 2014 ericqUpdateLSCNotch at 110MHz

Quote:

Q >> Please measure the RF spectrum again with the notch. 

The notch filter has been installed directly attached to the output of the SHP-150 at the PD output. Structurally, there is a right angle SMA elbow between the two filters; I set up a post holder under the notch pomona box to prevent torque on the PD. Via directional coupler and AG4395, we measured the output of the REFL165 RF amplifier with the PRMI locked on REFL33. 

Note, the plot below is not referred to the amplifier output, as in my previous plots; it is directly representative of the amplifier output spectrum. 

165_Notched.pdf

There are no RF signals being output above -28dBm, thus I am confident that we are not subject to compression distortion. 

Given the last measurements we made (ELOG 10692), I estimate that the notch has reduced the power at 110MHz by ~33dB, which is 9dB higher than the notch performance Koji measured when he made it. Maybe this could be due to the non-50Ohm impedance of the HPF distorting the tuning, or I physically detuned it when mounting it on the PD. Still, 33dB is pretty good, and may even give us room to amplify further. (ZRL-700+ instead of the ZFL-1000LN+?)

 

  10698   Tue Nov 11 21:41:09 2014 KojiUpdateLSC3f DRMI sensing mat

Sensing matrix calculation using DTT + Matlab

Note: If the signal phase is, for example,  '47 deg', the phase rotation angle is -47deg in order to bring this signal to 'I' phase.

Note2: As I didn't have the DQ channels for the actuation, only the relative signs between the PDs are used to produce the radar chart.
This means that it may contain 180deg uncertainty for a particular actuator. But this does not change the independence (or degeneracy) of the signals.



=== Sensing Matrix Report ===
Test time: 2014-11-11 08:14:00
Starting GPS Time: 1099728855.0
 

== PRCL ==
Actuation frequency: 621.13 Hz
Suspension (PRM) response at the act. freq.: 5.0803e-14/f^2 m/cnt
Actuation amplitude: 20.3948 cnt/rtHz
Actuation displacement: 1.0361e-12 m/rtHz
 
C1:LSC-AS55_I_ERR_DQ 4.20e+10
C1:LSC-AS55_Q_ERR_DQ -1.91e+11
==> AS55: 1.95e+11 [m/cnt] -24.58 [deg]
C1:LSC-REFL11_I_ERR_DQ 3.17e+12
C1:LSC-REFL11_Q_ERR_DQ -8.04e+10
==> REFL11: 3.17e+12 [m/cnt] -18.20 [deg]
C1:LSC-REFL33_I_ERR_DQ 4.15e+11
C1:LSC-REFL33_Q_ERR_DQ 4.28e+10
==> REFL33: 4.17e+11 [m/cnt] -137.11 [deg]
C1:LSC-REFL55_I_ERR_DQ 1.90e+10
C1:LSC-REFL55_Q_ERR_DQ -9.91e+09
==> REFL55: 2.14e+10 [m/cnt] -58.58 [deg]
C1:LSC-REFL165_I_ERR_DQ -1.16e+11
C1:LSC-REFL165_Q_ERR_DQ -3.14e+10
==> REFL165: 1.20e+11 [m/cnt] 45.20 [deg]
 
 
== MICH ==
Actuation frequency: 675.13 Hz
Suspension (ITMX) response at the act. freq.: 1.0312e-14/f^2 m/cnt
Suspension (ITMY) response at the act. freq.: 1.0224e-14/f^2 m/cnt
Actuation amplitude: 974.2957 cnt/rtHz
Actuation displacement (ITMX+ITMY): 2.0007e-11 m/rtHz
 
C1:LSC-AS55_I_ERR_DQ 2.55e+12
C1:LSC-AS55_Q_ERR_DQ 4.51e+12
==> AS55: 5.18e+12 [m/cnt] 113.51 [deg]
C1:LSC-REFL11_I_ERR_DQ -4.84e+10
C1:LSC-REFL11_Q_ERR_DQ -4.07e+09
==> REFL11: 4.85e+10 [m/cnt] 168.06 [deg]
C1:LSC-REFL33_I_ERR_DQ 2.06e+10
C1:LSC-REFL33_Q_ERR_DQ -9.39e+09
==> REFL33: 2.26e+10 [m/cnt] -167.51 [deg]
C1:LSC-REFL55_I_ERR_DQ 2.52e+09
C1:LSC-REFL55_Q_ERR_DQ -1.02e+10
==> REFL55: 1.05e+10 [m/cnt] -107.09 [deg]
C1:LSC-REFL165_I_ERR_DQ -1.79e+10
C1:LSC-REFL165_Q_ERR_DQ -5.50e+10
==> REFL165: 5.79e+10 [m/cnt] 102.02 [deg]



== SRCL ==

Actuation frequency: 585.13 Hz
Suspension (SRM) response at the act. freq.: 5.5494e-14/f^2 m/cnt
Actuation amplitude: 1176.3066 cnt/rtHz
Actuation displacement: 6.5278e-11 m/rtHz
 
C1:LSC-AS55_I_ERR_DQ -9.90e+10
C1:LSC-AS55_Q_ERR_DQ -1.18e+11
==> AS55: 1.54e+11 [m/cnt] -76.89 [deg]
C1:LSC-REFL11_I_ERR_DQ 2.96e+08
C1:LSC-REFL11_Q_ERR_DQ 4.78e+08
==> REFL11: 5.62e+08 [m/cnt] 41.42 [deg]
C1:LSC-REFL33_I_ERR_DQ -2.93e+09
C1:LSC-REFL33_Q_ERR_DQ 1.23e+10
==> REFL33: 1.27e+10 [m/cnt] -39.63 [deg]
C1:LSC-REFL55_I_ERR_DQ 3.71e+09
C1:LSC-REFL55_Q_ERR_DQ 2.78e+09
==> REFL55: 4.63e+09 [m/cnt] 5.86 [deg]
C1:LSC-REFL165_I_ERR_DQ -1.80e+10
C1:LSC-REFL165_Q_ERR_DQ 2.68e+10
==> REFL165: 3.23e+10 [m/cnt] -26.02 [deg]
 


Demodulation phases of the day

    'C1:LSC-AS55_PHASE_R = -53'
    'C1:LSC-REFL11_PHASE_R = 16.75'
    'C1:LSC-REFL33_PHASE_R = 143'
    'C1:LSC-REFL55_PHASE_R = 31'
    'C1:LSC-REFL165_PHASE_R = 150'

Attachment 1: DRMI_radar.pdf
DRMI_radar.pdf
  10697   Tue Nov 11 19:46:35 2014 diegoUpdateComputer Scripts / ProgramsStatus of the new nodus

The new nodus machine is being brought to life; until installation is finished and everything is fine, the old nodus will be unharmed. For future reference:

New Nodus hardware:

Case: SuperMicro SC825MTQ-R700U

M/B: SuperMicro X8DTU-F

CPU: 2x Intel Xeon X5650

RAM: 3x Kingston KVR1333D3S8R9S (2GB)

         3x Samsung M393B5673EH1-CH9 (2GB)

           Total 12 GB

HDD: Seagate ST3400832AS (400GB)

 

Current software situation and current issues :

1) Ubuntu Server 12.04.5 is installed and updated

2) The usual 'controls' user is present, with UID=1001 and GID=1001

3) Packages installed: nfs-common, rpcbind, ntp, dokuwiki, apache2, php5, openssh-server, elog-2.8.0-2 [from source], make, gcc, libssl-dev [dependencies for elog], subversion

4) Network: interface eth0 is set up (static IP and configuration in /etc/network/interfaces); eth1 is recognized and added, but not configured yet

5) DNS: configuration is in /etc/resolvconf/resolv.conf.d/base (since /etc/resolv.conf is overwritten by the resolvconf program using the 'base' database)

6) ntp is installed and (presumably) configured, but ntpd misbehaves (namely, all the servers are found, but a 

tail /var/log/syslog

shows that no actual synchronization is performed, and the daemon keeps

Listening on routing socket on fd #22 for interface updates

7) dokuwiki apache2 php subversion elog are installed but not configured yet (I need info about their current state, configuration and whereabouts)

8) I copied and merged the old nodus' .bashrc and .cshrc into new nodus' .bashrc, need to know if something has to be added

9) backup frames, backup user dirs and 40m public_html are not set yet, as in #7

 

Is there something missing?

If there is something missing from here (ligo/cds software, smartmontools/hddtemp and similar, or anything else) tell me and I'll set them up.

  10696   Tue Nov 11 03:48:46 2014 JenneUpdateLSC3f DRMI

I was able to lock the DRMI on 3f signals this evening, although the loops are not stable, and I can hear oscillations in the speakers.  I think the big key to making this work was the placement of the SHP-150 high pass filter at the REFL165 PD, and also the installation of Koji's 110 MHz notch filter just before the amplifier, which is before the demod board for REFL165.  These were done to prevent RF signal distortion.

DRMI 3f:   With DRMI locked on 1f (MICH gain = 1, PRCL gain = -0.05, SRCL gain = 2, MICH = 1*REFL55Q, PRCL = 0.1*REFL11I, SRCL = 1*REFL165I), I excited lines, and found the signs and values for 3f matrix elements.  I was using the same gains, but MICH = 0.5*REFL165Q, PRCL = 0.8*REFL33I and SRCL = -0.2*REFL165I.  Part of the problem is likely that I need to include off-diagonal elements in the input matrix to remove PRCL from the SRCL error signal. 

With the DRMI locked on 1f, I took a sensing matrix measurement.  I don't think we believe the W/ct of the photodiode calibration (we need to redo this), but otherwise the sensing matrix measurement should be accurate.  Since the calibration of the PDs isn't for sure, the relative magnitude for signals between PDs shouldn't be taken as gospel, but within a single PD they should be fine for comparison. 

As a side note, we weren't sure about the MICH -> ITMs balancing, so we checked during a MICH-only, and with the locking apparatus we are unable to measure a difference between 1's for both ITMs in the output matrix, and 1 for ITMX and 0.99 for ITMY.  So, we can't measure 1% misbalances in the actuator, but we think we're at least pretty close to driving pure MICH. 

We kind of expect that SRCL should only be present in the 55 and 165 PDs, although we see it strongly in all of the REFL PDs.  Also, PRCL and SRCL are not both lined up in the I-phase.  So, I invite other people to check what they think the sensing matrix looks like. 

  • The excitation lines (and matching notches) were on from 12:14am (
  • Nov 11 2014 08:14:00 UTC / GPS 1099728856) to 12:20am (
  •  
  • Nov 11 2014 08:20:00 UTC / GPS 
  • 1099729216) for 360sec. 
  • MICH was driven with 800 counts at 675.13 Hz, with +1*ITMY, -1*ITMX. 
  • PRCL was driven with 1000 counts at 621.13 Hz with the PRM. 
  • SRCL was driven with 800 counts at 585.13 Hz using the SRM. 

All 3 degrees of freedom have notches at all 3 of those frequencies in the FM10 of the filter banks (and they were all turned on).  During this time, DRMI was locked with 1f signals. 

DRMI sensing matrix:

 SensMatMeas_10Nov2014_DRMI.png

Earlier in the evening, I also took a PRMI sensing matrix, with the PRMI locked on REFL33 I&Q.  Watch out for the different placement of the plots - I couldn't measure AS55 in the DRMI case, since cdsutils.avg freaked out if I asked for more than 14 numbers (#PDs * #dofs) at a time.

SensMatMeas_10Nov2014_PRMI.png

Rana, Koji and I were staring at the DRMI sensing matrix for a little while, and we aren't sure why PRCL and SRCL aren't co-aligned, and why they aren't orthogonal to MICH.  Do we think it's possible to do something to digitally realign them?  Will the solution that we choose be valid for many CARM offsets, or do we have to change things every few steps (which we don't want to do)? 

Things to work on:

* Reanalyze DRMI sensing matrix data from 12:14-12:20am. 

* Make a simulated scan of higher order mode resonances in the arm cavities.  Is it possible that on one or both sides of the CARM resonance we are getting HOM resonances of the sidebands? 

* Figure out how to make DRMI 3f loops stable.

* Try CARM offset reduction with DRMI, and / or PRMI on REFL 165.

  10695   Tue Nov 11 01:38:23 2014 KojiUpdateLSCNotch at 110MHz

To further reduce the RF power at 110MHz in the REFL165 chain, I made a twin-t notch in a pomona box.

It is tuned at 110.66MHz.

The inductor is Coil Craft 5mm tunable (164-09A06SL 100-134nH).
Without the 10Ohm resister (like a usual notch), the dip was ~20dB. With this configuration, the notch of -42dB was realized.

Q >> Please measure the RF spectrum again with the notch.

 

Attachment 1: twin_t_notch.pdf
twin_t_notch.pdf
Attachment 2: notch_tf.pdf
notch_tf.pdf
  10694   Mon Nov 10 23:14:20 2014 ranaSummarySUSSRM: damping gains & Optical Lever servo Tune-up
  1.  tweaked gains for POS, PIT, YAW, SIDE by ~10-20% to get nice ringdowns with Q~5.
  2. Measured bounce rool freqs = 16.43 / 23.99 Hz. Updated the Mech Res Wiki page. Tightened up the bandstops to get back a few deg of phase. Propagated these new bandstops into the SRM SUS damping filter banks.
  3. Made and turned on a LP filters for the loops.
  4. Added a ~0.3 Hz gain boost / bubble.
  5. Set UGF to be ~2.5x below where it rings up. Estimate this to be ~3.5 Hz.
  6. First PDF shows bounce/roll peaks in OSEMs. Notice how f_roll is > sqrt(2)*f_bounce. ??
  7. Second PDF shows the OL spectra with the loops on/off. Previously there was no Boost and no LP (!!) turned on.
  8. 3rd PDF shows the modeled Bode plot of the OLG. yellow/blue is boost off/on
Attachment 1: SRM-BR.pdf
SRM-BR.pdf
Attachment 2: SRM_err_141110.pdf
SRM_err_141110.pdf
Attachment 3: SRM-OLG.pdf
SRM-OLG.pdf
  10693   Mon Nov 10 18:23:10 2014 ericqConfigurationLSCDRMI sensing

ARG, I accidentally permuted the digital demod angles. This significantly weakens the argument for believing AS55I is broken... In fact, Jenne and I did some investigations this afternoon that showed that the channel is indeed working. SRX error signal strangeness remains unexplained, however. 

Also, I have yet to compensate for the gain of the violin filters; the actuator calibration numbers I used were for the SUS-LSC FMs, not the LSC FMs where I was injecting. New measurements will be taken soon, as well, since REFL165 is hopefully improved. 

Corrected plots are below. 

REFL11.pngREFL33.png

REFL55.pngREFL165.png

AS55.png

  10692   Mon Nov 10 18:11:57 2014 ericqUpdateLSC3F RFPD RF spectra

 Jenne and I measured the situation using a SHP-150 directly attached to the REFL165 RF output, and at first glance, the magnitude of the 165MHz signal seems to not be distorted by the amplifier. 

 165signals.pdf

We will soon investigate whether 165 signal quality has indeed improved. 

  10691   Mon Nov 10 07:36:35 2014 SteveUpdateSUSPRM damping recovered

PRM sus damping recovered and PMC locked.

  10690   Sat Nov 8 16:01:32 2014 ericqConfigurationLSCDRMI sensing

Here are some preliminary results from the sensing sweeps I did the other night. 

Notes:

  • The analog AS55I signal chain is almost certainly busted in some way. This would also explain the odd looking error signals in SRX, and was actually hypothesized by Koji when discussing the SRX oddness. 
  • I used the same mirror calibration numbers from Koji's recent Elogs to turn these into counts/m.
  • MICH was excited via differential ITM motion.  I also performed a TF with BS driven MICH, with the compensating PRM output matrix in place, and it looks different, but I haven't looked too deeply into it yet. 
  • The angles plotted are in regard to the analog I and Q signals (i.e., I took TFs to I_ERR and Q_ERR and then unrotated by the digital rotation angle); this is why I suspect AS55I is broken, as all of the signals are entirely in the analog Q.
  • The amplitudes seem to be roughly consistent with Koji's recent observations. 
  • I still need to cut out the violin-filter-corrupted data points to quote the sensing elements with error bars...

Plots!

 REFL11.pngREFL33.png

REFL55.pngREFL165.png

AS55.png

xml files, and DttData matlab script used to generate these plots is attached. 

Attachment 6: DRMIsensing.zip
  10689   Sat Nov 8 11:35:05 2014 ranaUpdateLSC3F RFPD RF spectra

 

 I think 'saturation' here is a misleading term to think about. In the RF amplifiers, there is always saturation. What we're trying to minimize is the amount of distorted waveforms appearing at 3f and 15f from the other large peaks. Usually for saturation we are worried about how much the big peak is getting distorted; not the case for us.

  10688   Sat Nov 8 11:31:51 2014 ranaUpdateComputer Scripts / ProgramsAnaconda + CDSutils

Quote:

I've fallen down the rabbit hole of trying to reconcile our desire for newer versions of the Numpy and Scipy python packages with the use of our handy cdsutils tools. 

 Avoid rabbit holes! What I did at LLO which works is to install an Anaconda in some shared directory and then just make an alias which puts that directory at the head of the path when running the more advanced SciPy installs. It works fine and cannot interfere with our usual operation since its only sourced when running peak find.

  10687   Fri Nov 7 17:44:10 2014 diegoUpdateLSCIR Resonance Script Status

Yesterday I did some more tests with a modifies script; the main difference is that scipy's default wavelet implementation is quite rigid, and it allows only very few choices on the wavelet. The main issue is that our signal is a real, always positive symmetrical signal, while wavelets are defined as 0-integral functions, and can be both real or complex, depending on the wavelet; I found a different wavelet implementation, and I combined it with some modified code from the scipy source, in order to be able to select different wavelets. The result is the wavelet_custom.py module, which lives in the same ALS script directory and it is called by the script. In both the script and the module there the references I used while writing them. It is now possible to select almost any wavelet included in this custom module; "almost" means that the scipy code that calls the find_peaks_cwt routine is picky on the input parameters of the wavelet function, I may dig into that later. For the last tests, instead of using a Ricker wavelet (aka Mexican hat, or Derivative of Gaussian Order 2), I used a DOG(6), as it also has two lesser positive lobes, which can help in finding the resonance; the presence of negative lobes is, as I said, unavoidable. I attach an example of the wavelet forms that are possible, and in my opinion, excluding the asymmetric and/or complex ones, the DOG(6) seems the best choice, and it has provided slightly better results. There are other wavelet around, but they are not included in the module so I should implement them myself, I will first see if they seem fitting our case before starting writing them into the module. However, the problem of not finding the perfect working point (the "overshoot-like" plot in my previous elog) is not completely solved. Eric had a good idea about that: during the fine scan, the the PO*11_ERR_DQ signals should be in their linear range, so I could also use them and check their zero crossing to find the optimal working. I will be working on that.

Attachment 1: wavelets.nb.zip
  10686   Fri Nov 7 16:15:53 2014 JenneUpdateLSC3F RFPD RF spectra

I have found an SHP-150, but no SHP-175's (also, several 200's, and a couple of 500's).

Why do you say the SHP-150 isn't enough?  The blue peak at 10*fmod in your plot looks like it's at -12 dBm.  -13 dB on top of that will leave that peak at -25 dBm.  That should be enough to keep us from saturation, right?  It's not a lot of headroom, but we can give it a twirl until a 175-er comes in.  

Koji also suggests putting in a 110 MHz notch, combined with either an SHP-150 or SHP-175, although we'll have to measure the combined TF to make sure the notch doesn't spoil the high pass's response too much.

Quote:

165_HPatPD.png


Minicircuits' SHP-150 only has 13dB suppression at 110MHz, which would not be enough either. SHP-175 has 31dB suppression at 110MHz and 0.82dB at 160MHz, maybe this is what we want.

 

  10685   Fri Nov 7 14:41:18 2014 ericqUpdateLSC3F RFPD RF spectra

Quote:

 After some enlightening conversation with Koji, we figured that the RF amplifier in the REFL165 chain is probably being saturated.

The measurements I took yesterday bear this out. However, even putting the high-pass directly on the PD output doesn't reduce the signal enough to avoid saturating the amplifier.

We need to think of the right way to get the 165MHz signal at large enough, but undistorted, amplitude to the demod board. 


 The current signal chain looks like:

AS Table                                  LSC RACK
[ PD ]----------------------------------->[ AMP ]------>[ 100MHzHPF ]----->[ DEMOD ]
      (1)                                        (2)                 (3)

I previously made measurements at (3). Let's ignore that. 

Last night, I took measurements with a directional coupler at points (1) and (2), to see the signal levels before and after the amplifier. I divided the spectrum at (2) by the nominal gain of the amplifier, 23.5dB; thus if everything was linear, the spectra would be very similar. This is not the case, and it is evident why. There are multiple signals stronger than -20dBm, and the amplifier has a 1dB compression point of +3dBm, so any one of these lines at 4x, 6x and 10x fMod is enough to saturate. 

 165_ampSaturation.png


I also made a measurement at point 4 in the following arrangement, in an attempt to reduce the signal amplitude incident on the amplifier.  

AS Table                                           LSC RACK
[ PD ]->[ 100MHzHPF ]----------------------------------->[ AMP ]--------->[ DEMOD ] 
                                                                (4) 

 Though the signals below 100MHz are attenuated as expected, the signal at 110MHz is still too large for the amplifier. 

165_HPatPD.png


Minicircuits' SHP-150 only has 13dB suppression at 110MHz, which would not be enough either. SHP-175 has 31dB suppression at 110MHz and 0.82dB at 160MHz, maybe this is what we want.

  10684   Fri Nov 7 10:29:56 2014 SteveUpdateGeneralHe/Ne telescope

Quote:

Liyuan is measuring the He/Ne telescopes in the Y arm between the tube and CES wall. He'll be here till 1pm

 Liyuan is continuing his measurement in the Y arm till noon today.

  10683   Fri Nov 7 02:21:12 2014 ericqUpdateLSC3F RFPD RF spectra

 

After some enlightening conversation with Koji, we figured that the RF amplifier in the REFL165 chain is probably being saturated (the amp's 1dB compression is at +3dBm, has 23dB gain, and there are multiple lines above -20dBm coming out of the PD). I took a few more spectrum measurements to quantify the consequences, as well as a test with the highpass connected directly to the PD output, that should reduce the power into the amplifier. However, I am leaving everything hooked back up in its original state (and have removed all couplers and analyzers...)

I also took some DRMI sensing measurements. In the simple Michelson configuration, I took TFs of each ITMs motion to AS55Q to make sure the drives were well balanced. They were. Then, in the DRMI, I took swept sine TFs of PRCL, SRCL and differential ITM MICH motion to the Is and Qs of AS55 and all of the REFLs. I constrained the sweeps to 300Hz->2kHz; the loops have some amount of coupling so I wanted to stay out of their bandwidth. I also took a TF of the pure BS motion and BS-PRM MICH to the PDs. From these and future measurements, I hope to pursue better estimates of the sensing matrix elements of the DRMI DoFs, and perhaps the coefficients for compensating both SRCL and PRCL out of BS motion. 

I'm leaving analysis and interpretation for the daytime, and handing the IFO back to Diego...

ELOG V3.1.3-